Suppose that Congress sets the top personal tax rate on interest and dividends at 38% and the top rate on realized capital gains at 17%. The corporate tax rate stays at 23%. Assume capital gains are half of equity income.


a. Compute the difference between the total corporate plus personal taxes paid on debt and the total taxes on equity income if all capital gains are realized immediately. (Do not round intermediate calculations. Round your answer to 4 decimal places.)


b. Compute the difference between the total corporate plus personal taxes paid on debt and the total taxes on equity income if all capital gains are deferred forever. (Do not round intermediate calculations. Round your answer to 4 decimal places.)
Answer #1

Part A

For debt income:

Corporate tax = $0

Personal tax = 0.380

Total = $0.380

For equity income, with all capital gains realized immediately:

Corporate tax = 0.230

Personal tax = (0.38*0.5*(1-0.23)) + (0.17*0.5*(1-0.23)) = 0.2118

Total = $0.4418

Part B

Corporate tax = 0.230

Personal tax = 0.38*0.5*(1-0.23) = 0.1463

Total = $0.3763